Ashley makes deposits of $293 monthly (at the end of each month) into an account that pays 1.7% interest compounded monthly.

How much money will be in the account after 5 years?


1
Expert's answer
2021-04-15T07:25:37-0400

Given the following:

P=$293, r= 1.7%(0.017), n= 5, and m= 12

The FV of the deposits is given as "FV=\frac{A{(1+\frac{r}{m})^{mn}}}{\frac{r}{m}}"

"FV=\frac{293{(1+\frac{0
.017}{12})^{12×5}}}{\frac{0.017}{12}}"

"FV=183661.93"


The amount in the account after 5 years will be $183661.93